A solution of hydrochloric acid of a concentration of 3.6 g/dm3 is titrated against 25 cm3 of sodium hydroxide. What is the molarity of the solution of hydrochloric acid?


1
Expert's answer
2021-03-22T06:37:40-0400

Volume of sodium hydroxide = 25/1000 = 0.0250dm3


Amount of sodium hydroxide = 1×0.00250 = 0.00250mol


NaOH + HCl > NaCl + H20


mole ratio NaOH : HCL is 1:1


Hence 0.00259 mole of NaOH react with 0.00250 mole of HCl


 concentration of HCL = 3.6g/dm3/ 36.5 ( relative formula.


= 0.136 mol/dm3 


volume of HCl = 0.00250/0.136


= 0.0184 dm3


concentration of HCL = 0.0025 mol/0.0184


= 0.1359mol/dm3